OpenStudy (anonymous):

A pendulum has a mass of 3 kg and is lifted to a height of 0.3 m. What is the maximum speed of the pendulum?

OpenStudy (anonymous):

is the equation 1/2mv^2 ? i thought so but i dont have a "v" so im not sure

OpenStudy (lastdaywork):

Does this question comes with a diagram??

OpenStudy (anonymous):

no @LastDayWork

OpenStudy (lastdaywork):

Use energy balance - PE of initial case = KE of final case

OpenStudy (anonymous):

what do you mean? @LastDayWork

OpenStudy (lastdaywork):

1/2mv^2 = mgh

OpenStudy (lastdaywork):

I should have written it ^^ in reverse Change in Potential energy = Change in Kinetic energy mgh = 1/2mv^2 Actually, there is a - sign in the above equation. But personally i will suggest to find only magnitude from mathematics and try to deduce sign by using physics.

OpenStudy (jfraser):

use the conservation of energy, as @LastDayWork suggests.

OpenStudy (anonymous):

what should i use for v? @LastDayWork

OpenStudy (lastdaywork):

@abbyisaprincess Srry, i didn't check the notification. The question asks you to find the max speed. This is what you will get by energy balance.

OpenStudy (jfraser):

@abbyisaprincess the v is the only thing you don't know, so rearrange the equation to solve for the v

OpenStudy (anonymous):

1/2 x m x v^2 = m x g x h cancel out the m and solve for v. therefore v^2= (g x h) / 0.5 = square root of [(9.8 x 0.3) / 0.5]